OpenStudy (anonymous):

a 10 Hz water wave travels from deep water, where its speed is 38 cm/s, to shallow water, where its speed is 28 cm/s and angle1=30. Find the index of refraction.

OpenStudy (anonymous):

any formulas?

OpenStudy (anonymous):

v1/v2=L1/L2 n1sinA1=n2sinA2 sinA1/sinA2=v1/v2=L1/L2=n2/n1

OpenStudy (anonymous):

L=wavelength n1= index of refraction of first medium n2=index of refraction of second medium

OpenStudy (anonymous):

thats not in my book

OpenStudy (anonymous):

how will I find the different light speeds